# Availability-aware Sensor Fusion via Unified Canonical Space for 4D Radar, LiDAR, and Camera

### v1

@@ C.1 Additional Qualitative Results on Damaged LiDAR
Table 5: Performance comparison of ASF under various sensor combinations evaluated on the K-Radar [12] benchmark v2.0. We indicate the employed sensors (C: Camera, L: LiDAR, R: 4D Radar) and report $AP_{3D}$ at IoU = 0.3 under various weather conditions for the ‘Sedan’ and ‘Bus or Truck’ classes. The bold and underlined values indicate the best and the second-best, respectively. All ten ASF models share the same neural network weights trained with R+L+C.
@@
Table 6: Performance comparison of ASF under various sensor combinations evaluated on the K-Radar [12] benchmark v2.0. We indicate the employed sensors (C: Camera, L: LiDAR, R: 4D Radar) and report $AP_{BEV}$ at IoU = 0.3 under various weather conditions for the ‘Sedan’ and ‘Bus or Truck’ classes. The bold and underlined values indicate the best and the second-best, respectively. All ten ASF models share the same neural network weights trained with R+L+C.
@@
Table 7: Performance comparison of ASF under various sensor combinations evaluated on the K-Radar [12] benchmark v2.0. We indicate the employed sensors (C: Camera, L: LiDAR, R: 4D Radar) and report $AP_{3D}$ at IoU = 0.5 under various weather conditions for the ‘Sedan’ and ‘Bus or Truck’ classes. The bold and underlined values indicate the best and the second-best, respectively. All ten ASF models share the same neural network weights trained with R+L+C.
@@
Table 8: Performance comparison of ASF under various sensor combinations evaluated on the K-Radar [12] benchmark v2.0. We indicate the employed sensors (C: Camera, L: LiDAR, R: 4D Radar) and report $AP_{BEV}$ at IoU = 0.5 under various weather conditions for the ‘Sedan’ and ‘Bus or Truck’ classes. The bold and underlined values indicate the best and the second-best, respectively. All ten ASF models share the same neural network weights trained with R+L+C.
C.2 Additional Quantitative Results
Tables 5, 6, 7, and 8 present comprehensive evaluations of ASF performance across various sensor configurations and weather conditions. Results demonstrate that multi-sensor fusion approaches consistently outperform single-sensor methods. The C+L+R configuration achieves the highest performance in most scenarios for both ‘Sedan’ and ‘Bus or Truck’ classes, with C*+L+R often ranking second, which indicates that a damaged camera does not significantly affect performance.
Performance analysis across weather conditions reveals that radar-inclusive combinations (C+L+R, L+R) exhibit superior performance in adverse weather (demonstrating 4D Radar’s effectiveness under challenging environmental conditions), particularly in fog and snow. While camera-LiDAR fusion (C+L) occasionally outperforms in normal and overcast conditions, its performance deteriorates significantly in precipitation.

### v2

@@ D.2 Additional Quantitative Results
Table 10: Performance comparison of ASF under various sensor combinations evaluated on the K-Radar (kradar) benchmark v2.0. We indicate the employed sensors (C: Camera, L: LiDAR, R: 4D Radar) and report $AP_{BEV}$ at IoU = 0.3 under various weather conditions for the ‘Sedan’ and ‘Bus or Truck’ classes. The bold and underlined values indicate the best and the second-best, respectively. All ten ASF models share the same neural network weights trained with R+L+C.
@@
Table 11: Performance comparison of ASF under various sensor combinations evaluated on the K-Radar (kradar) benchmark v2.0. We indicate the employed sensors (C: Camera, L: LiDAR, R: 4D Radar) and report $AP_{3D}$ at IoU = 0.5 under various weather conditions for the ‘Sedan’ and ‘Bus or Truck’ classes. The bold and underlined values indicate the best and the second-best, respectively. All ten ASF models share the same neural network weights trained with R+L+C.
@@
Table 12: Performance comparison of ASF under various sensor combinations evaluated on the K-Radar (kradar) benchmark v2.0. We indicate the employed sensors (C: Camera, L: LiDAR, R: 4D Radar) and report $AP_{BEV}$ at IoU = 0.5 under various weather conditions for the ‘Sedan’ and ‘Bus or Truck’ classes. The bold and underlined values indicate the best and the second-best, respectively. All ten ASF models share the same neural network weights trained with R+L+C.
@@
Acknowledgment
This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. 2021R1A2C3008370).